Print

Print


 Ticket Number: HEA-NOC/20020802-01         Ticket Status: CLOSE
 Ticket Type:   Scheduled                   Ticket Source: HEA-NOC
 Ticket Scope:  Site                        Site/Line:     HEAnet <==> NCIR
 Ticket Owner:  [log in to unmask]   Problem Fixer: HEA-NOC
 Ticket Opened:                             Problem Start: 20020802 12:15:00 UTC+1
 Ticket Closed: 20020807  09:08:21 UTC      Problem End:   20020807 09:47 UTC+1


Problem Description:

NCI notify that their circuit and services to are go offline as part of preparation for migration to new premises over the course of the weekend.

Affected:

All traffic to and from NCI. MX backup is at at Kerna.

Actions:



Time to Fix:

 Additional Notes:

Move is complete, Links to both existing and new campuses are active